# Simple Guidance Mechanisms for Discrete Diffusion Models

### v1

@@ Appendix B Relating UDLM to CTMC
B.1 Equivalence of UDLM ELBO and SEDD ELBO
Below, we demonstrate that the variational lower bound from (18) used to train UDLM is equivalent to the lower bound derived in SEDD (Lou et al., 2023).
@@ B.1 Equivalence of UDLM ELBO and SEDD ELBO
From (1) we have $q_{t}({\mathbf{z}}|{\mathbf{x}})=\alpha_{t}{\mathbf{x}}_{i}+(1-\alpha_{t})/N$ where $i$ is the non-zero index of the one-hot vector ${\mathbf{z}}$, i.e., ${\mathbf{z}}_{i}=1$. Thus, the ‘true’ conditional score in (B.1) can be written as
@@
In our work, we use the mean parameterization: we predict the ‘clean’ data given noisy observations using the model that we denote as ${\mathbf{x}}_{\theta}$. Note that (B.1) is minimized if
@@
We can now use the identities from (B.1) and (51) to demonstrate that $\text{NELBO}_{\text{SEDD}}$ (B.1) is equivalent to $\text{NELBO}_{\text{UDLM}}$ (18).
@@
Appendix C Experimental Details
C.1 Dataset Details
@@
The QM9 dataset comes from Ruddigkeit et al. (2012) and Ramakrishnan et al. (2014). The dataset is comprised of $\sim$133k small molecules. We process the dataset using the RDKit library (Landrum et al., 2013) to extract ‘canonical’ SMILES string representations and add the annotations for drug-likeness (QED) and ring-counts. The data are tokenized using a regular expression from Schwaller et al. (2019). Input lengths of 32 tokens were used for pre-training and generation experiments.
@@
This dataset was downloaded from https://huggingface.co/datasets/billion-word-benchmark/lm1b. We tokenize using the bert-base-uncased tokenizer. Sequences were padded to a max input length of 128 tokens.x
@@ C.3 Training Configurations
C.4 Guidance Details
C.4.1 Baselines
@@
In our QM9 experiments, $topk$ is set as the vocabulary size, which is 40. When training the classifier, we use a smaller classification model using the same dataset specified in Appendix C.1. The smaller backbone DiT consists of 8 layers, 8 attention heads, and a hidden dimension of 512. We use the same hyperparameters as those specified in Table 14.
@@
For the QM9 experiments, when training the classifier for PPLM, we use the same dataset specified in Appendix C.1. We use the same hyperparameters as those specified in Table 14, except the LR peak is reduced to $3\mathrm{e}^{-5}$ and minimum is reduced to $3\mathrm{e}^{-7}$.
@@
To implement the NOS baseline from Gruver et al. (2024), we train a classifier where we use the same backbone as the unconditional diffusion model (see Appendix C.2 for architecture details), and we initialize and freeze weights using the pre-trained unconditional diffusion model. We then mean pool the last hidden embeddings of the backbone and linearly project them to the classification logits. This final projection layer represents the only trainable parameters of the guidance model.
@@
For the QM9 experiments, when training the classifier for NOS, we use the same dataset specified in Appendix C.1. We use the same hyperparameters as those specified in Table 14, except the LR peak is reduced to $3\mathrm{e}^{-5}$ and minimum is reduced to $3\mathrm{e}^{-7}$.
C.4.2 D-CFG Details
@@
C.4.3 D-CBG Details
For D-CBG in the QM9 experiment, we train a smaller classification model using the same dataset specified in Appendix C.1. The smaller backbone DiT consists of 8 layers, 8 attention heads, and a hidden dimension of 512. We apply mean pooling on the final hidden representations before linearly projecting to the classification logits. We use the same hyperparameters as those specified in Table 14.
@@ C.4.3 D-CBG Details
C.5 Guidance Evaluation Details
C.5.1 Genomic Sequences Metrics
@@
To compute the $k$-mer distirbution shift, for each species, we create counts for each of the unique 3mers and 6mers for that species in the validation set and 64 generated sequences for that species. We then compute the Jensen-Shannon divergence between those categorical histograms. Finally we take a weighted average of these distances across species where the weights are given by the relative species proportion in the validation dataset, see Table 11 for the relative proportions.
@@
Finally, controllability is measured by the macro-averaged F1 of an oracle model on the 640 generated sequences. Our oracle model is a separate HyenaDNA model with 8 layers and hidden dimension 256, which has 6.6M parameters. This model was downloaded and initialized from scratch from https://huggingface.co/LongSafari/hyenadna-small-32k-seqlen-hf. We mean pool the final layer embeddings and linearly project them to the ten category classification logits. This model was trained on the full Species10 dataset as described in Appendix C.1. For reference, in Table 15 we present the classification results of the this oracle model in the 5% valdiation set of the original data. We see that other than difficulty distinguishing between human and gorilla genomes, the model can serve as a near perfect oracle.
@@
C.5.2 CIFAR10 Quality Metrics
@@ C.5.3 QM9 Guidance Metrics
Appendix D Guidance Ablation Results
D.1 Effect of Varying Guidance Hyperparameters
@@
$\gamma$ $T$ 3-mer JS ($\downarrow$) 6-mer JS ($\downarrow$) Disc. AUROC ($\downarrow$) F1 ($\uparrow$)
@@
D.2 Effect of Using First Order Approximation in D-CBG
The data for this analysis is available in Tables 21, 24, 30, and 33. We present it graphically in Figure 6. In these plots, we use $T=32$ for diffusion generation.
@@ D.3 Effect of Varying $T$
Appendix E Unconditional Sample Generation

### v3

@@ Appendix C Relating UDLM to CTMC
C.1 Equivalence of UDLM ELBO and SEDD ELBO
Below, we demonstrate that the variational lower bound from (18) used to train UDLM is equivalent to the lower bound derived in SEDD (Lou et al., 2023), for a specific parameterization of the denoising score matching.
@@ C.1 Equivalence of UDLM ELBO and SEDD ELBO
From (1) we have $q_{t}({\mathbf{z}}|{\mathbf{x}})=\alpha_{t}{\mathbf{x}}_{i}+(1-\alpha_{t})/N$ where $i$ is the non-zero index of the one-hot vector ${\mathbf{z}}$, i.e., ${\mathbf{z}}_{i}=1$. Thus, the ‘true’ conditional score in (C.1) can be written as
@@
In our work, we use the mean parameterization: we predict the ‘clean’ data given noisy observations using the model that we denote as ${\mathbf{x}}_{\theta}$. Note that (C.1) is minimized if
@@
We can now use the identities from (C.1) and (56) to demonstrate that $\text{NELBO}_{\text{SEDD}}$ (C.1) is equivalent to $\text{NELBO}_{\text{UDLM}}$ (18).
@@
Appendix D Experimental Details
D.1 Dataset Details
@@
The QM9 dataset comes from Ruddigkeit et al. (2012) and Ramakrishnan et al. (2014). The dataset is comprised of $\sim$133k small molecules. We process the dataset using the RDKit library (Landrum et al., 2013) to extract ‘canonical’ SMILES string representations and add the annotations for drug-likeness (QED) and ring-counts. The data are tokenized using a regular expression from Schwaller et al. (2019), which is available here: https://huggingface.co/yairschiff/qm9-tokenizer. We used input lengths of 32 tokens (with right-sided padding) for pre-training and generation experiments.
@@
This dataset was downloaded from https://huggingface.co/datasets/billion-word-benchmark/lm1b. We tokenize using the bert-base-uncased tokenizer. Sequences were padded to a max input length of 128 tokens.
@@ D.3 Training Configurations
D.4 Guidance Details
D.4.1 Baselines
@@
In our QM9 experiments, $topk$ is set as the vocabulary size, which is 40. When training the classifier, we use a smaller classification model using the same dataset specified in Appendix D.1. The smaller backbone DiT consists of 8 layers, 8 attention heads, and a hidden dimension of 512. We use the same hyperparameters as those specified in Table 13.
@@
For the QM9 experiments, when training the classifier for PPLM, we use the same dataset specified in Appendix D.1. We use the same hyperparameters as those specified in Table 13, except the LR peak is reduced to $3\mathrm{e}^{-5}$ and minimum is reduced to $3\mathrm{e}^{-7}$.
@@
To implement the NOS baseline from Gruver et al. (2024), we train a classifier where we use the same backbone as the unconditional diffusion model (see Appendix D.2 for architecture details), and we initialize and freeze weights using the pre-trained unconditional diffusion model. We then mean pool the last hidden embeddings of the backbone and linearly project them to the classification logits. This final projection layer represents the only trainable parameters of the guidance model.
@@
For the QM9 experiments, when training the classifier for NOS, we use the same dataset specified in Appendix D.1. We use the same hyperparameters as those specified in Table 13, except the LR peak is reduced to $3\mathrm{e}^{-5}$ and minimum is reduced to $3\mathrm{e}^{-7}$.
D.4.2 D-CFG Details
@@
D.4.3 D-CBG Details
For D-CBG in the QM9 experiment, we train a smaller classification model using the same dataset specified in Appendix D.1. The smaller backbone DiT consists of 8 layers, 8 attention heads, and a hidden dimension of 512. We apply mean pooling on the final hidden representations before linearly projecting to the classification logits. We use the same hyperparameters as those specified in Table 13.
@@ D.4.3 D-CBG Details
D.5 Guidance Evaluation Details
D.5.1 Genomic Sequences Metrics
@@
To compute the $k$-mer distirbution shift, for each species, we create counts for each of the unique 3mers and 6mers for that species in the validation set and 64 generated sequences for that species. We then compute the Jensen-Shannon divergence between those categorical histograms. Finally we take a weighted average of these distances across species where the weights are given by the relative species proportion in the validation dataset. See Table 10 for the relative proportions.
@@
Finally, controllability is measured by the macro-averaged F1 of an oracle model on the 640 generated sequences. Our oracle model is a separate HyenaDNA model with 8 layers and hidden dimension 256, which has 6.6M parameters. This model was downloaded and initialized from scratch from https://huggingface.co/LongSafari/hyenadna-small-32k-seqlen-hf. We mean pool the final layer embeddings and linearly project them to the ten category classification logits. This model was trained on the full Species10 dataset as described in Appendix D.1. For reference, in Table 14 we present the classification results of the this oracle model on the 5% valdiation set of the original data. We see that other than difficulty distinguishing between human and gorilla genomes, the model can serve as a near perfect oracle.
@@
D.5.2 CIFAR10 Quality Metrics
@@ D.5.3 QM9 Guidance Metrics
Appendix E Guidance Ablation Results
E.1 Effect of Varying Guidance Hyperparameters
@@
D-CFGγ $T$ 3-mer JS ($\downarrow$) 6-mer JS ($\downarrow$) Disc. AUROC ($\downarrow$) F1 ($\uparrow$)
@@
E.2 Effect of Using First Order Approximation in D-CBG
The data for this analysis is available in Tables 20, 23, 29, and 32. We present it graphically in Figure 5. In these plots, we use $T=32$ for diffusion generation.
@@ E.3 Effect of Varying $T$
Appendix F Unconditional Sample Generation